i have prepared hydroxyapatite from calcium hydroxide and orthophosphoric acid. Can there be a carbonate peak in FTIR analysis? If so where did carbonate come from?
1
Expert's answer
2017-05-08T08:28:09-0400
Calcium hydroxide could absorb some CO2 from the air: Ca(OH)2 + CO2 = Ca(HCO3)2 or Ca(OH)2 + CO2 = CaCO3 + H2O. However, under H3PO4 action it (CaCO3 calcium carbonate) cannot survive. Thus, there will be no CO32- anion. Check carefully your background for CO2 which is always there.